# Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming

### v1

@@ 5 Discussion and Conclusion
Evidence that ICL is implicitly a gradient-based learning mechanism
We started with the question of whether ICL could be a processing mechanism of LLMs that can capture the flexibility and adaptability of human learning mechanisms. One type of learning mechanism that supports flexible adaptation to data is gradient descent. Motivated by previous theoretical proposals that ICL could involve functional gradient descent or implicit fine-tuning, we attempted to better characterize what kind of learning ICL is in actual models. We focused on one particular aspect of ICL: whether it involves an implicit gradient component during the forward computation. We differ from previous approaches for analyzing ICL in that we test standard LLMs with natural language data. Having established a connection between ICL and human structural priming, we used the IFE to diagnose whether LLMs behave as if they are implicitly performing gradient descent when processing a prime sentence. We found that LLMs do indeed display the IFE in many cases, particularly in larger models. These findings support the hypothesis that a gradient computation is implicitly performed in the forward pass of ICL. Differences between models suggests that ICL as a gradient-descent-like process arises only in larger models. More speculatively, these results raise the possibility that gradient-based learning might be a crucial property that enables generalization from a small number of samples, an ability that is shared by human learners and (to some extent) LLMs performing ICL. Our study not only provides behavioral results that align LLMs’ behaviors with human behaviors on structural priming at the processing mechanism level, but also demonstrates the possibility of studying the nature of ICL with off-the-shelf pre-trained LLMs and with naturalistic data.
ICL as a consequence of language modeling?
ICL is typically understood as involving demonstration-answer pairs in the prompt. Inspired by data-centric views that explain ICL from the distributional properties of pre-training data, we proposed a generalized notion of ICL that is sensitive to general parallelisms. As a result, any text in the prompt could serve as an implicitly defined “task” of following the template provided in the context and generating a parallel structure. Therefore, ICL could be viewed as a side product of the general propensity for structural parallelism in the language modeling task. We leave this perspective for future investigation.
@@
In this study, we focused on one single “task”, namely selecting DO versus PD sentence structures. If our reasoning is correct, such that it is indeed the implicit gradient component of the ICL that results in LLMs’ capability of capturing the IFE, then we predict that the IFE diagnostics could be generalized to other ICL tasks, even non-linguistic tasks. Future work could extend our current method to additional ICL tasks. Finding the IFE on a wider range of tasks would strengthen the claim that ICL is driven by implicit gradient computation, while not observing the IFE on other tasks would indicate that ICL only sometimes displays a gradient-descent-like character.
@@
Although ICL is generally identified as a phenomenon at the behavioral level, having an explanation at the mechanistic level is desirable since it brings greater interpretability and contributes to more concrete theory building. Our current study, despite using real pre-trained models and naturalistic data, remains at the behavioral level and is empirical in nature. Given our current contribution of establishing a connection between ICL and human priming and using the IFE as a way to diagnose whether ICL works in a gradient-descent-like manner or not, future work could improve our understanding by incorporating techniques from mechanistic interpretability to explain our current finding at the mechanistic level.
@@
As ICL capability is often argued to scale with model size, we predict in Section 4.2 that the IFE effect will be more robust in larger models. Although the difference in the IFE behavior between GPT2 and GPT3-base is substantial, we do not believe that we have observed a saturation of the IFE because the $R^{2}$ values that have observed are far from the ceiling level of 1.0. GPT3-base is currently the biggest model on which we have access to the logit predictions, but we believe the same behavioral test could be applied to larger models in order to verify our prediction.
@@ Acknowledgments
We would like to thank the anonymous reviewers from the ESSLLI Student Session 2024 and AMLaP 2024 for their thoughtful comments, which helped us refine this paper. We would also like to thank the members of the Yale Linguistics Department for helpful discussion. Any errors remain our own.
@@ Appendix A Finding pronoun probabilities in the OpenWebText corpus with spaCy
As is mentioned in Section 3.2, we constructed the WithPronoun version of the corpus in order to investigate the impact of animacy of the indirect object on the verb biases. To do this, we approximated the distribution of the natural occurrence frequencies over the set of English pronouns in dative alternation sentences from a fragment of the OpenWebText corpus (Gokaslan and Cohen, 2019), which is a reasonable proxy for the closed WebText corpus that was used to train the GPT2 models.
We parsed a fragment (around 160 million tokens) of the corpus with spaCy (Honnibal et al., 2020). Specifically, we used the en_core_web_trf specification of the spaCy model, and we identified the set of dative alternation sentences by doing dependency parsing on each sentence. Then, we counted the frequencies of the set of English pronouns that occurred as the indirect object of the ditransitive verb. The list of pronouns and their frequencies are presented in Table 2, sorted by frequency.

### v2

@@ 5 Discussion and Conclusion
Evidence that ICL is implicitly an error-driven learning mechanism
We started with the question of whether ICL could be a processing mechanism of LLMs that can capture the flexibility and adaptability of human learning mechanisms. One type of error-driven learning mechanism that supports flexible adaptation to data is gradient descent. Motivated by previous theoretical proposals that ICL could in principle be implicitly performing gradient descent or fine-tuning, we attempted to better characterize what kind of learning ICL is in actual models. We focused on one particular aspect of ICL: whether it involves an implicit error signal during the forward computation. In contrast to previous analyses of ICL, we test standard LLMs with natural language data.
Having established a connection between ICL and human structural priming, we used the IFE to diagnose whether LLMs behave as if they are implicitly performing error-driven learning when processing a prime sentence. The very nature of the IFE requires an error-driven mechanism: processing must in some way be sensitive to the error signal, i.e., the amount of divergence from a model’s expectations. In the case of DO/PD sentences, the IFE is defined based on verb biases: the error signal is created by the degree of mismatch between the expectation on structural alternatives based on the verb bias of the prime verb and the actual perceived prime structure. The psycholinguistic literature discussed in Section 2.1 was the first to identify the logic of this connection between the IFE and error-driven learning, and it is this logic that we are employing in the framing of this paper. Note however that our argumentation does not in any way assume that ICL involves the same mechanism as human implicit learning.
We found that LLMs do indeed display the IFE in many cases, particularly in larger models. These findings support the hypothesis that an error signal is implicitly computed in the forward pass of ICL. Differences between models suggests that ICL only takes on a gradient-descent-like nature in larger models. More speculatively, these results raise the possibility that error-driven learning might be a crucial property that enables generalization from a small number of samples, an ability that is shared by human learners and (to some extent) LLMs performing ICL. Our study not only provides behavioral results that align LLM and human behavior in structural priming at the processing mechanism level, but also demonstrates the possibility of studying the nature of ICL with off-the-shelf pre-trained LLMs and with naturalistic data.
ICL as a consequence of language modeling
ICL is typically understood as involving demonstration-answer pairs in the prompt. Inspired by data-centric views that explain ICL from the distributional properties of pre-training data , we proposed a generalized notion of ICL that is sensitive to general parallelisms. Therefore, ICL could be viewed as a side product of the general propensity for structural parallelism in the language modeling task. We leave this perspective for future study.
@@
In this study, we focused on one single “task”, namely selecting DO versus PD sentence structures. If our reasoning is correct, such that it is indeed the implicit error signal of the ICL that results in LLMs’ capability of capturing the IFE, then we predict that the IFE diagnostics could be generalized to other ICL tasks, even non-linguistic tasks. Future work could extend our current method to additional ICL tasks. Finding the IFE on a wider range of tasks would strengthen the claim that ICL is driven by implicit error-driven computation, while not observing the IFE on other tasks would indicate that ICL only sometimes displays a gradient-descent-like character.
@@
Although ICL is generally identified as a phenomenon at the behavioral level, having an explanation at the mechanistic level would bring greater interpretability and could contribute to more concrete theory building. Our current study, despite using real pre-trained models and naturalistic data, remains at the behavioral level and is empirical in nature. Given our current contribution of establishing a connection between ICL and human priming and using the IFE as a way to diagnose whether ICL works in a gradient-descent-like manner or not, future work could improve our understanding by incorporating techniques from mechanistic interpretability to explain our current finding at the mechanistic level.
@@
As ICL capability is often argued to scale with model size, we predict in Section 4.2 that the IFE effect will be more robust in larger models. Although the difference in the IFE behavior between GPT2 and GPT3-base is substantial, we do not believe that we have observed a saturation of the IFE because increasing the confidence level to $99\%$ could still result in confidence intervals containing 00. GPT3-base is currently the biggest model on which we have access to the logit predictions, but we believe the same behavioral test could be applied to larger models in order to verify our prediction.
@@ Acknowledgments
We would like to thank the anonymous reviewers from the ARR June and October 2024 cycles, ESSLLI Student Session 2024, and AMLaP 2024 for their thoughtful comments, which helped us refine this paper. We would also like to thank the members of the Computational Linguistics at Yale Lab, the Yale Computation and Cognition Joint Lab, and the Yale Linguistics Department for helpful feedback. We thank the Yale Center for Research Computing for guidance and use of the research computing infrastructure, specifically the Grace cluster. Any errors are our own.
@@ Appendix A Finding pronoun probabilities in the OpenWebText corpus with spaCy
As is mentioned in Section 3.2, in order to estimate the verb biases represented in the training corpus of GPT2 models, we parsed a fragment (approximately 160 million tokens) of the OpenWebText corpus (Gokaslan and Cohen, 2019) with the Python package spaCy (Honnibal et al., 2020) to get an estimate of the DO vs. PD ratio for each verb. Specifically, we used the en_core_web_trf specification of the spaCy model, and we identified the set of dative alternation sentences by doing dependency parsing on each sentence. We found that the verb biases from the corpus are less well-represented in GPT2 models, motivating our decision to estimate verb biases by using model judgments rather than corpus statistics; see Section 3.4.
The corpus parse results motivated us to investigate the impact of whether the indirect object is a pronoun on the verb biases. Thus, we constructed the WithPronoun version of the corpus. To do this, we approximated the distribution of the natural occurrence frequencies over the set of English pronouns in dative alternation sentences from a fragment of the OpenWebText corpus (Gokaslan and Cohen, 2019), which is a reasonable proxy for the closed WebText corpus that was used to train the GPT2 models. Then, we counted the frequencies of the set of English pronouns that occurred as the indirect object of the ditransitive verb. The list of pronouns and their frequencies are presented in Table 2, sorted by frequency.
@@ Appendix B Fine-tuning details
Appendix C Verb Biases represented in LLMs
See Figure 5 for details of the verb biases in GPT3.
Figure 5: Comparison of PD biases with (left) and without (right) pronoun for GPT3. As is shown in Equation. 1, a high PD-bias means a larger proportion of probability assigned to the PD structure against the DO structure in LLMs.
Appendix D Full Statistics of the IFE Experiments in Concatenation Mode
See Table 4 (as a supplement to Table 1) for the full statistical metrics of the IFE experiments reported in Section 4.2.
Std_Err $r$-value $p$-value $R^{2}$ RMSE Model Pronoun Obj? PD-PD DO-PD PD-PD DO-PD PD-PD DO-PD PD-PD DO-PD PD-PD DO-PD GPT2-small True 0.021 0.018 0.119 -0.088 0.597 0.698 0.014 0.008 0.020 0.017 GPT2-small False 0.020 0.023 0.154 0.058 0.495 0.798 0.024 0.003 0.016 0.019 GPT2-medium True 0.023 0.017 -0.123 -0.327 0.585 0.138 0.015 0.107 0.023 0.016 GPT2-medium False 0.019 0.031 -0.258 -0.245 0.247 0.273 0.067 0.060 0.017 0.027 GPT2-large True 0.019 0.018 0.129 -0.416 0.568 0.054 0.017 0.173 0.019 0.018 GPT2-large False 0.020 0.027 -0.038 -0.163 0.868 0.469 0.001 0.026 0.018 0.024 Llama-7b True 0.016 0.014 -0.270 -0.803 0.224 0.000 0.073 0.645 0.015 0.013 Llama-7b False 0.027 0.059 -0.215 -0.387 0.336 0.076 0.046 0.149 0.019 0.042 Llama-7b-chat True 0.019 0.018 -0.138 -0.766 0.541 0.000 0.019 0.587 0.018 0.017 Llama-7b-chat False 0.036 0.066 -0.082 -0.327 0.718 0.137 0.007 0.107 0.024 0.044 Llama-13b True 0.019 0.012 -0.568 -0.872 0.006 0.000 0.323 0.760 0.018 0.011 Llama-13b False 0.034 0.074 -0.400 -0.473 0.065 0.026 0.160 0.224 0.019 0.042 GPT3-davinci-002 True 0.015 0.013 -0.755 -0.813 0.000 0.000 0.570 0.662 0.013 0.011 GPT3-davinci-002 False 0.031 0.055 -0.415 -0.507 0.055 0.016 0.172 0.257 0.020 0.035
Table 4: The standard error, $r$-value, $p$-value, $R^{2}$ coefficient, and RMSE score for each condition under the Concatenation mode.